# Big Tech influence over AI research revisited: memetic analysis of attribution of ideas to affiliation

### v1

@@ 4.2 Most contagious ideas in AI research
Using the meme score measure, we were able to identify and analyze which ideas are the most prevalent in AI research in general, which is the first step in differentiating Big Tech from Academia.
@@
To perform this analysis, we calculated the meme scores, and propagation scores and counted the occurrences for all memes present in our dataset, which resulted in over 1.7 million observations. Such a vast amount of data had to be filtered to get the most interesting results, that is why we decided to limit our study to the phrases where the meme score was greater than 0, which resulted in over 60 000 meaningful observations.
@@
Since our main goal is to deeply analyze the most spreading ideas, we limited the number of memes according to two criteria: the observed occurrences and their meme score values. The first criterion is based on the idea that the memes that have high meme scores but only occur once or twice are not really meaningful topics, as they are too niche to be considered contagious. The second condition is obvious, since the higher the meme score, the more likely it is to spread. To choose a cutoff value, we analyzed the curves shown in Fig. 5. Finally, after limiting ourselves to phrases with a meme score above 0.25 and at least 20 occurrences, we ended up with 251 observations representing the top memes in our dataset.
@@
The set of chosen top memes was later annotated by the group of domain experts which is briefly described in C, resulting in the assignment of one category per observation. The resulting categories are ML algorithms, Medical terms, Security, Niche ML applications, ML Concept, Maths, Graphs, Computer Vision, Natural Language Processing, Data Related, and Other. The annotations of the top memes allow us to analyze which high-level topics are popular among ML researchers, indicating the most promising research areas. According to Table 1, the most popular topic among scientists is the medical applications of ML. This category has the most memes and a third number of the memes occurrences in our dataset. It is not the only application, however, as we can see on the top memes list other important topics like security, and a whole category describing niche applications of ML solutions. However, this doesn’t mean that today’s research focuses only on ML applications, as the second-highest average meme score value, and the second sum of the occurrences belong to the topic describing the ML algorithms. Additionally, we were able to distinguish some particularly popular directions of AI development, such as Computer Vision or NLP. Finally, such an in-depth analysis proves the quality of our framework as the category Other, which includes unrelated topics, as well as meaningless memes, is responsible for only 20% of detected memes. The group has the biggest number of occurrences due to the memes such as ‘paper’, ‘results’, or ‘art’ (part of ‘state of the art’ phrase), which occur multiple times in the majority of the papers.
@@
As we already know the big picture of the main topics in the scientist’s interests, we can combine it with a more detailed analysis, and provide a wider perspective for our top memes group.
Figure 6 shows the characteristics of the top memes subset. As the meme score value depends on the number of memes occurrences as well as the propagation score we can easily imagine that the line that stands for the meme score equal to 0.25 is used as a cutoff value in this analysis. From the density of observations, especially in our group, we can grasp the tendency that the highest meme scores were assigned for the observations with a smaller number of occurrences. Additionally, the plot helps us detect the outliers aforementioned in the previous section, as they yield low propagation score, but also have a large number of occurrences. Thanks to the exemplary meme annotations we can also track some memes and analyze the reasons why are they considered important. For example, ‘fairness’ is considered important mostly due to its high propagation score, whereas ‘cnn’ or ‘deep learning’ benefits from a vast number of occurrences.
To finalize the topic analysis, we decided to present all top memes in the form of a structured word cloud presented in Fig. 7. This way, we are able to analyze particular memes that build each of the predefined categories. We should particularly focus on the medical memes, as this topic is definitely the most important one, due to the abundance of different memes. The memes creating this topic (presented as orange), cover various medical terms. We can discover which disease treatments and studies greatly benefit from the usage of ML (ex. various cancers, epilepsy, diabetic retinopathy, Alzheimer’s disease, malaria, etc.). We can also say that another major points of interest are the studies considering genomics (lncRNA, miRNAs, microRNAs, etc.), and that ML is widely used for various examinations (mammograms, eeg, ecg, wireless capsule endoscopy, etc.).
@@ 4.3 Differences in contagiousness between Companies, Academia, and Big Tech
In order to perform a more systematical analysis, we compared all non-zero conditioned sticking factors across all affiliation groups using Kruskal-Wallis test. Results indicate a significant difference across groups in terms of the sticking factor, i.e. affiliation is a factor that creates distinctions among sticking factor values. Outcomes of post-hoc comparison performed using Dunn’s test between categories presented in Fig. 8b pinpoint distinctions among the sticking factor distributions: strikingly, there are no statistical differences among Academia, Big Tech, and Company. On the other hand, such pairs as mixed Big Tech – Big Tech or mixed Company – Company can be distinguished. This is exactly opposite to the relations observed using network metrics for non-isolated nodes and shown in D, where differences between mixed Big Tech – Big Tech and mixed Company – Company were negligible while one could clearly distinguish Academia, Big Tech, and Company categories.
The above analysis is conducted in absolute terms, i.e., the distributions can be biased by the fact that Academia affiliations are exceedingly more frequent than Big Tech or Company ones (cf Fig. 2). To overcome this issue we focused on memes that appear in at least two different affiliation categories, e.g., “random neural network” (present in mixed Company, Company, mixed Big Tech, and Big Tech) or “captcha” (present in Company, mixed Big Tech, and Big Tech) depicted in Fig. 8a. Following, we conducted Wilcoxon signed-rank tests (Wilcoxon, 1945), comparing sticking factors conditioned on specific groups using matched samples (the same memes in both groups). In case we were able to reject the null hypothesis (observations in both groups and are exchangeable) in favor of the one-sided alternative hypothesis (differences are stochastically larger than a distribution symmetric about zero) we noted it with a directed link in Fig. 8c, pointing from to . The results revealed that memes affiliated with Big Tech exhibited higher sticking factors compared to all other affiliation categories, with company-affiliated memes ranking second. Academic affiliation, on the other hand, was associated with a decrease in the sticking factor compared to other affiliations.
Following this, we explore how this difference depends on the meme category introduced in Table 1. Five hundred memes with the highest meme scores, excluding those not present at least once in papers from each affiliation group, were selected for this comparison. After removing duplicates and the least common categories of memes, the resulting list has 148 memes. The resulting comparison of conditioned sticking factors divided by meme category is presented in Fig. 9.
@@ 5 Discussion and conclusions
This work aimed to identify the most contagious ideas (memes) in AI research and assess to what extent the contagiousness of memes depends on the affiliation group of authors of the paper, and whether this contagiousness differs between pure and mixed academic or company affiliation. For this purpose, we analyzed AI papers’ abstracts, affiliations, and citation networks. We employed two key metrics, namely the meme score and the sticking factor, to quantify the contagiousness of these memes. Additionally, we introduced the concept of conditioned sticking factor to assess the relationship between affiliations and contagiousness.
@@
Our findings reveal several noteworthy insights. Firstly, we observed that papers affiliated with Big Tech tend to receive disproportionate citation counts in comparison to all companies. However, the most highly cited papers are those co-authored by individuals from both Big Tech and academic institutions, i.e., papers having mixed affiliations. Additionally, papers co-authored by individuals from both corporate and academic backgrounds exhibit scientometric distinctions from papers authored exclusively by Academia or corporate entities. This pattern holds true for both Big Tech and Company, thus challenging the binary distinctions often assumed in prior research.
@@
The analysis of contagiousness indicates that on average, the contagiousness of Academia-authored memes does not differ from from those authored by companies or Big Tech, although when taking into account only memes present in both affiliations, the same memes have on average higher contagiousness when mentioned by Big Tech than Academia. These facts together could be linked with the difference in the diversity of memes produced and propagated by those two groups. Academia creates far more memes than Big Tech or Companies (mostly due to the difference in overall papers authored), resulting in mean contagiousness being not significantly different than Big Tech or Company. When comparing common ideas of interest of those two groups, the difference in contagiousness becomes apparent.
The impact of affiliation group over the contagiousness of memes could be also understood by applying the framing theory. Framing theory states that an issue can be perceived from various perspectives, each carrying implications for multiple values or considerations (Chong and Druckman, 2007). By applying this theoretical lens to the realm of memetics, we could view issues as memes. In this context, we assert that each meme possesses the potential for diverse framings. These framings, tied to the affiliations of the authors, become integral components that propagate alongside the meme. Consequently, the contagiousness of a meme, when conditioned on a specific group, signifies the extent of that group’s influence over the framing of the respective meme. In essence, the more contagious a meme becomes within a particular group, the more pronounced is that group’s control over the framing associated with the meme. To assess this further, one could analyze the co-occurrence of memes, and ties of such memes “conglomerates” to affiliations.
Further direction of research should also take into account the dynamics of memes contagiousness, investigating whether the impact of affiliation on contagiousness – potentially representing a gap between Academia, Companies, and Big Tech – is expanding or contracting would provide valuable insights.
A limitation of this study, inherent in the original meme score calculating technique, is the operationalization of memes as simple noun chunks, overlooking the potential synonymity of phrases. Consequently, there is a possibility that high-spreading memes may be overlooked. Possible solutions could be techniques such as clustering.
The analysis which we have presented confronts the influence of Big Tech with that of academic institutions on the propagation of ideas between scientific papers. In doing so, we offer a unique perspective on tracing the evolution of ideas influenced by both major technological companies and academic institutions. The conditioned sticking factor, introduced in this work, represents a novel tool that can extend beyond the confines of this study. It can be employed to analyze how various modalities, such as country of origin, journal, or institution, impact the spread of ideas in different contexts.
@@
In conclusion, our results suggest that the prevalent notion of Big Tech’s dominance over AI research is overly simplistic. We advocate for a more nuanced understanding of the roles played by Big Tech companies, corporations, and Academia in shaping the future of AI. We believe that this nuanced perspective, as presented in our work, will contribute to more informed and reasoned discussions about the trajectory of AI research.
5.0.1 Acknowledgements
@@
The research was funded by (POB Cybersecurity and Data Science) of Warsaw University of Technology within the Excellence Initiative: Research University (IDUB) programme. This work was also funded by the European Union under the Horizon Europe grant OMINO – Overcoming Multilevel INformation Overload (grant number 101086321, http://ominoproject.eu). Views and opinions expressed are those of the authors alone and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor the European Research Executive Agency can be held responsible for them. Computational part of this study was supported in part by the Poznań Supercomputing and Networking Center (grant number 607).
@@ Appendix A Data gathering and processing
Following Liu et al. [2021] we employed simple search strategy for AI papers. We searched for following words from titles and abstracts: artificial intelligence, machine learning, classifier, neural network, deep learning, data science, nlp, machine-learning, computer vision.
@@ Appendix C Annotations
The annotation process was conducted for the top memes, defined as the ones that have a meme score value over 0.25, and more than 20 occurrences, by the team of domain experts. The pipeline for the creation of the annotations consisted of several steps, described below.
@@
Appendix D Network analysis without isolated nodes
Here we perform the same analysis as carried out in Sec. 4.1, however, we restrict the original dataset by removing zero-degree nodes. Regardless of the type of the last category (i.e., Big Tech or company) as well as of the variable in question (in-degree, PageRank) Kruskal-Wallis test finds significant differences among all groups (p-value ). However, as can be seen in Figs 11 and 12 post-hoc tests indicate a lack of difference between mixed category and Big Tech (or company) one, underlining the importance of zero-degree nodes in this context.

### v2

@@ 4.2 Most contagious ideas in AI research
Using the meme score measure, we identified and analyzed which ideas are the most prevalent in AI research in general, which is the first step in differentiating Big Tech from Academia. To perform this analysis, we calculated the meme scores and propagation scores and counted the occurrences for all memes in our dataset, resulting in over 1.7 million observations. Such an amount of data had to be filtered to get the most interesting results. Thus, we decided to limit our study to phrases where the meme score was greater than 0, resulting in over 60,000 meaningful observations.
@@
Since our main goal is to analyze the most spreading ideas deeply, we limited the number of memes according to two criteria: the observed occurrences and their meme score values. The first criterion is based on the idea that the memes with high meme scores that only occur once or twice are not meaningful topics, as they are too niche to be considered contagious. The second condition is obvious since the higher the meme score, the more likely it is to spread. We analyzed the curves shown in Fig. 5 to choose a cutoff value. Finally, after limiting ourselves to phrases with a meme score above 0.25 and at least 20 occurrences, we ended up with 251 observations representing the top memes in our dataset. A manual annotation and inspection of the selected memes yields 80% precision, which is a comparable result to the one reported by Kuhn et al. 2014 (i.e., around 81.2% on 150 memes). Let us underline that it is possible to exclude obviously meaningless memes like ‘paper’ or ‘results,’ however, it requires introducing a free parameter into Eq. (1), in a similar way as in Kuhn et al. (2014). Another option is a manual intervention. We refrained from following either of these paths to maintain high clarity of methodology.
@@
The set of chosen top memes was later annotated by the group of domain experts, which is briefly described in C, resulting in the assignment of one category per observation. The resulting categories are ML algorithms, Medical terms, Security, Niche ML applications, ML Concept, Maths, Graphs, Computer Vision, Natural Language Processing, Data Related, and Other. The annotations of the top memes allow us to analyze which high-level topics are popular among ML researchers, indicating the most promising research areas. According to Table 1, the most popular topic among scientists is the medical applications of ML. This category has the most memes and a third number of the memes occurrences in our dataset. It is not the only application, however, as we can see on the top memes list other important topics like security, and a whole category describing niche applications of ML solutions. However, this does not mean that today’s research focuses only on ML applications, as the second-highest average meme score value, and the second sum of the occurrences belong to the topic describing the ML algorithms. Additionally, we were able to distinguish some particularly popular directions of AI development, such as Computer Vision or NLP. Finally, such an in-depth analysis proves the quality of our framework as the category Other, which includes unrelated topics, as well as meaningless memes, is responsible for only 20% of detected memes. The group has the biggest number of occurrences due to the memes such as ‘paper’, ‘results’, or ‘art’ (part of ‘state of the art’ phrase), which occur multiple times in the majority of the papers.
@@
Knowing the big picture of scientist’s interests, we can combine it with a more detailed analysis, providing a wider perspective for our top memes group.
Figure 6 shows the characteristics of the top memes subset. As the meme score value depends on the number of meme occurrences and the propagation score, we can imagine that reflecting a meme score equal to 0.25 is used as a cutoff value in this analysis. From the density of observations, we can grasp the tendency for the highest meme scores to be assigned to the observations with fewer occurrences. Additionally, the plot helps us detect the outliers aforementioned in the previous section, as they yield low propagation score, but also have many occurrences. Thanks to the exemplary meme annotations, we can also track some memes and analyze the reasons why they are considered important. For example, ‘fairness’ is considered important mostly due to its high propagation score, whereas ‘cnn’ or ‘deep learning’ benefits from a vast number of occurrences.
To finalize the topic analysis, we present all top memes as a structured word cloud shown in Fig. 7. This way, we can analyze particular memes that build each of the predefined categories. We should particularly focus on the medical memes, as this topic is definitely the most important one due to the abundance of different memes. The memes creating this topic (presented as orange) cover various medical terms. We can discover which disease treatments and studies greatly benefit from the usage of ML (e.g., various cancers, epilepsy, malaria, etc.). We also can indicate that another major points of interest are the studies considering genomics (lncRNA, miRNAs, etc.), and that ML is widely used for various examinations (mammograms, eeg, ecg, etc.).
@@ 4.3 Differences in contagiousness between Companies, Academia, and Big Tech
In a more systematical analysis, we compared all non-zero conditioned sticking factors across all affiliation groups using Kruskal-Wallis test, indicating a significant difference across groups in terms of the sticking factor, i.e., affiliation is a factor that creates distinctions among sticking factor values. Outcomes of post-hoc comparison (Dunn’s test) between categories presented in Fig. 8b pinpoint distinctions among the sticking factor distributions: strikingly, there are no statistical differences among Academia, Big Tech, and Company. On the other hand, such pairs as mixed Big Tech – Big Tech or mixed Company – Company can be distinguished. This is opposite to the relations observed using network metrics for non-isolated nodes shown in Figs. 3 and 4.
The above analysis is conducted in absolute terms, i.e., the distributions can be biased because Academia affiliations are more frequent than Big Tech or Company ones (cf Fig. 2). To overcome it, we focused on memes that appear in at least two different affiliation categories, e.g., “random neural network” (present in mixed Company, Company, mixed Big Tech, and Big Tech) depicted in Fig. 8a. Following, we conducted Wilcoxon signed-rank tests, comparing sticking factors conditioned on specific groups using matched samples (the same memes in both groups). If we were able to reject the null hypothesis (observations in both groups and are exchangeable) in favor of the one-sided alternative hypothesis (differences are stochastically larger than a distribution symmetric about zero), we noted it with a link in Fig. 8c, pointing from to . The results revealed that memes affiliated with Big Tech exhibited higher sticking factors compared to all other affiliation categories, with company-affiliated memes ranking second. On the other hand, academic affiliation was associated with a decrease in the sticking factor compared to other affiliations.
We then explored how this difference depends on the meme category introduced in Table 1. Five hundred memes with the highest meme scores, excluding those not present at least once in papers from each affiliation group, were selected for this comparison. After removing duplicates and the least common categories of memes, the resulting list has 148 memes. The comparison of conditioned sticking factors divided by meme category is presented in Fig. 9.
@@ 5 Discussion and conclusions
This study aims to identify and measure the contagiousness of AI research ideas (memes) based on authors’ affiliations and whether this differs between pure and mixed academic or company affiliations. We used two metrics, the meme score, and the sticking factor, and introduced the conditioned sticking factor to assess this relationship.
@@
Our research reveals that papers co-authored by individuals from both Big Tech and academic institutions tend to receive significantly higher citation counts. Additionally, papers co-authored by individuals from both corporate and academic backgrounds exhibit distinct scientific characteristics compared to papers authored exclusively by academia or corporate entities. These findings challenge the binary distinctions often assumed in prior research.
@@
The analysis suggests that while the contagiousness of memes authored by academia does not significantly differ from those by companies or Big Tech on average, memes mentioned by Big Tech have higher contagiousness when compared to academia. This difference in contagiousness could be linked to the diversity of memes produced and propagated by the two groups, with academia creating far more memes due to the difference in overall papers authored. The variance in contagiousness may also be attributed to factors such as the availability of resources for Big Tech’s researchers.
The impact of affiliation groups on the spread of memes can be understood through framing theory. This theory suggests that different perspectives and values influence how issues are perceived (Chong and Druckman, 2007). In the context of memes, the affiliations of creators contribute to different framings of the same meme. The contagiousness of a meme within a specific group indicates that group’s influence over the framing of the meme. Analyzing how memes co-occur and their connections to affiliations can provide further insight.
In our analysis, we explore the impact of Big Tech and academic institutions on the propagation of ideas in scientific papers. We introduce a new concept, the conditioned sticking factor, which can be used to analyze how various factors influence the spread of ideas. Our approach could help track relations between organizational characteristics and academia-industry interactions (see, e.g., Scandura and Iammarino, 2022) and possible outcomes could complement or challenge the results obtained in previous studies that tackled the problem of diffusion of ideas in a given scientific discipline (Hargreaves Heap and Parikh, 2005), or technology adoption under different (also institutional) constraints (Galang, 2014) Additionally, the conditioned meme score may aid in measuring information overload in science (Hołyst et al., 2024), a critical issue for researchers.
Further direction of research should consider the dynamics of memes’ contagiousness and investigate the impact of affiliation on contagiousness. Comparing the relationship between Big Tech and Academia in AI with similar relations in non-AI computer science and in other disciplines like medicine (Giunta et al., 2016) would provide valuable insights. A limitation of this study is the operationalization of memes as simple noun chunks, overlooking the synonymity of phrases and, consequently, potentially high-spreading memes. Possible solutions could be techniques such as clustering embedded memes. Another potential constraint comes from the method of selecting the keywords to filter AI-related papers – as mentioned in A, we only partially used the approach suggested by Liu et al. (2021), deeming another set of secondary and ternary keywords being suitable. One needs to mention that Färber and Tampakis (2023), who pursued a similar topic of academia and company affiliations in AI research, focused on the Microsoft Academic Graph field of study. In the same manner, Gao et al. (2024) relied on Web of Science classification, examining if patent-cited papers have a higher impact. All these examples show a lack of a common gold standard that might limit the generality of obtained conclusions. In recent years, non-profit research organizations like OpenAI have made notable contributions to the AI community, but their research output has not yet reached the scale of Big Tech (see Fig. 11 in D for specific numbers). These organizations, such as Anthropic, have innovative approaches to disseminating research findings using non-traditional publishing methods with interactive features. However, quantitatively analyzing their impact poses challenges, as their publications are excluded from standard paper indexing databases. Our analysis is also restricted to the year 2020, preventing the observation of memes related to generative AI applications.
@@
Despite these limitations, our study suggests that the notion of Big Tech’s dominance over AI research is overly simplistic, advocating for a more nuanced understanding of the roles played by Big Tech companies, corporations, and Academia in shaping the future of AI.
Acknowledgements
@@
The research was funded by (POB Cybersecurity and Data Science) of Warsaw University of Technology within the Excellence Initiative: Research University (IDUB) programme. This work was also funded by the European Union under the Horizon Europe grant OMINO – Overcoming Multilevel INformation Overload (grant number 101086321, https://ominoproject.eu). Views and opinions expressed are those of the authors alone and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor the European Research Executive Agency can be held responsible for them. Computational part of this study was supported in part by the Poznań Supercomputing and Networking Center (grant number 607).
@@ Appendix A Data gathering and processing
The article filtering strategy of our study is based on the work Liu et al. (2021). Its authors designed an AI-oriented search query that aims to cover various AI-related topics while maintaining high precision in extracted papers. To achieve these goals, they followed a three-step procedure:
1.
Retrieving AI benchmark records. The authors downloaded two narrow groups of papers, first from Web of Science (WoS), which include the ’artificial intelligence’ term as the topic of the papers, and the second which comes from the 19 top-tier AI-related journals. This way, the relevance of benchmark observations is ensured.
2.
Retrieving candidate keywords. Based on the benchmark records, the authors retrieve both ’Author Keywords’ and ’Keywords Plus’ to create a wide range of AI-related candidate keywords. Additionally, they remove the general terms (i.e. ’system’) and ascertain their meaning by checking online web sources like Wikipedia. This way, the list of 214 candidate keywords is created.
3.
Creating 3 keyword groups based on their relevance. Finally, the authors refine the candidate keywords by using co‑occurrence analysis and Hit Ratio screening.
The first group, called ’core’, contains ten terms that yield the highest co-occurrence numbers with the ’artificial intelligence’ keyword. The list includes the following terms: artificial intelligence, neural network, machine learning, expert system, natural language processing, deep learning, reinforcement learning, learning algorithms, supervised learning, intelligent agents.
In order to validate the remaining 204 terms, the authors calculated the Hit Ratio value for each term, which signifies how many records captured by a candidate keyword are also captured by the ’core’ lexical query. If the Hit Ratio was higher than 70%, then the keyword was added to the 2nd group; if it was between 30% and 70%, then it was manually checked if it was used in AI-related paper, and the score below 30% automatically excluded the candidate keyword.
To ensure the high relevance of AI-related papers in our study, we decided to follow the first filtering strategy, which includes 10 keywords. Additionally, we decided to include a subset of 19 terms that occur in the second and third groups and were deemed significant for the team of three domain (machine learning) experts. This additional list includes the following terms unsupervised learning, backpropagation learning, backpropagation algorithm, long short term memory, autoencoder, q learning, feedforward net, xgboost, transfer learning, gradient boosting, generative adversarial network, representation learning, random forest, support vector machine, multiclass classification, robot learning, graph learning, naive bayes classification, classification algorithm. Those filters were used in the first step of the data-gathering pipeline shown in Figure 10.
@@ Appendix C Annotations
The annotation process was conducted for the top memes, defined as the ones that have a meme score value over 0.25, and more than 20 occurrences, by three of domain experts. The pipeline for the creation of the annotations consisted of several steps, described below.
@@
Appendix D Google, Microsoft, and OpenAI citation count comparison